the motion of a particle is describe by x=10sin2t + 8cos2t. determine the period, amplitude and phase angle
1
Expert's answer
2019-05-21T10:47:02-0400
"x = 10 \sin 2t + 8 \cos 2t""x = 2 \sqrt{4^2 + 5^2} \bigg(\frac{5}{\sqrt{4^2 + 5^2}} \sin 2t + \frac{4}{\sqrt{4^2 + 5^2}} \cos 2t\bigg)"

Let us denote

"\cos \phi = \frac{5}{\sqrt{4^2+5^2}}"

then

"\sin \phi = \sqrt{1-\cos^2 \phi} = \sqrt{1 - \frac{5^2}{4^2 - 5^2}} = \frac{4}{\sqrt{4^2 + 5^2}}"


and

"\phi = \arctan \frac{4}{5}"

Using

"\cos \phi \sin 2t + \sin \phi \cos 2t = \sin (2t+\phi)"

we obtain

"x = 2 \sqrt{41} \sin(2t+\phi)"

The amplitude is

"2 \sqrt{41}"

the period is

"T = \frac{2\pi}{\omega} = \frac{2\pi}{2} = \pi"

and the phase angle is


"\phi = \arctan \frac{4}{5} = 0.67 \approx 39^\circ"